please help solve the following is a fraction

90 - 3/2n(fraction)=30

Answer by apachemexican1(64)

90-3/2n=30
subtract 90 from both sides of the equation
90-90-3/2n=30-90
-3/2n=-60
divide both sides by -3/2
n=40 good luck!